A block is chosen at random from a bag containing 6 white blocks, 4 black blocks, and 12 red blocks. What is the probability that it will be a red block or a white block?
Answer by stanbon(75887)

A block is chosen at random from a bag containing 6 white blocks, 4 black blocks, and 12 red blocks. What is the probability that it will be a red block or a white block?
---------------------
P(red or white) = P(red)+P(white) = (12/22)+(6/22) = 18/22
=====================================
Cheers,
Stan H.